OpenStudy (fools101):

sorry is there a Web Site i can do this is the same as the other one! y = 6x3 - 10x2 - 4x

OpenStudy (fools101):

please help!

OpenStudy (anonymous):

2x(3x^2-5x-2) 2x=0,x=0 (3x^2-5x-2)=(3x+1)(x-2)=0, x=2,x=-1/3

OpenStudy (fools101):

thank u here other!! y = x4 - 13x2 + 36 y = (x2 - 9)(x2 - 4)

OpenStudy (fools101):

please help!

OpenStudy (anonymous):

when you have (x^2-a) the factored pairs will be (x+sqrta)(x-sqrta)

OpenStudy (fools101):

so i would do it in both sides?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

doesnt look so scary now does it?

OpenStudy (fools101):

kind of so for this one i would be y = x4 - 13x2 + 36 y = (x2 - 9)/X2+9 = (x2 - 4)/ X2+4 like that?

OpenStudy (anonymous):

? you have (x^2-9)(x^2-4)=(x-3)(x+3)(x-2)(x+2)

OpenStudy (fools101):

ohhh !! kk can u show me example like something similar like that like how did u brake it down please! u don't have do it, just asking if u can just really need to understand this math and thank u for ur help! : )

OpenStudy (anonymous):

take x^2-25 25=a (x-sqrt(a))=(x-5) x+sqrt(a)=(x+5)

OpenStudy (anonymous):

anyway. I have to go. I have a class to teach in 20 min.

OpenStudy (fools101):

AWWWW kk!! thank u Soo Much!

OpenStudy (anonymous):

welcome.

OpenStudy (fools101):

^ ^have a nice Day!